OpenStudy (anonymous):

the line 1 passes through the points (9, -4) and has gradient 1/3. find an equation for line 1 in the form ax+by+c=0, where a,b and c are integers -the line l2 passes through the origin O and has gradient -2. the lines l1 and l2 intersect at the point P. calculate the coordinates of P.

OpenStudy (anonymous):

y-(-4)=(1/3)(x-9) y+4=(1/3)x-3 (1/3)x-y-7=0 or x-3y-21=0

OpenStudy (anonymous):

Thanks a lot :)

OpenStudy (anonymous):

any other doubts.....

OpenStudy (anonymous):

lol yaaa... the line l2 passes through the origin O and has gradient -2. the lines l1 and l2 intersect at the point P. calculate the coordinates of P.??

OpenStudy (anonymous):

which line 12.......?means

OpenStudy (anonymous):

line 2